OpenStudy (anonymous):

Jamie purchased a compact disk and paid $17.45. If the disk price is $15.99, what is the sales tax rate (round to the nearest tenth of a percent)? What's the tax rate?

OpenStudy (anonymous):

15.99+x=17.45 .. ?

OpenStudy (xapproachesinfinity):

think of it like this \(\large \rm 15.99+(15.99.a)=17.45\) if you paid 4$ for a purchase with sales tax of 7percent that mean you will pay 4+4x0.07=4.28$ in general if the sales tax is x or a or call it what ever if you the disk is 15.99 with the sales that it will be that amount add to it 15.99xa

OpenStudy (xapproachesinfinity):

you will find a decimal number multiply it by 100 to get your percentage

OpenStudy (xapproachesinfinity):

and don't forget to round as they asked

OpenStudy (xapproachesinfinity):

? did you find the answer

OpenStudy (anonymous):

i got 1.46%

OpenStudy (xapproachesinfinity):

wrong

OpenStudy (xapproachesinfinity):

how did you get that

OpenStudy (xapproachesinfinity):

i see you used a wrong equation that's why you got it wrong i said use this one 15.99+15.99x=17.45 solve for x to find your decimal number and then transfer it to percentage by multiplying by 100

OpenStudy (xapproachesinfinity):

I bet you didn't read my comment?

OpenStudy (xapproachesinfinity):

Do you know how to solve this kind of equation 15.99+15.99x=17.45

OpenStudy (anonymous):

I got 31.98 now i added 15.99+15.99

OpenStudy (xapproachesinfinity):

what? do you know how to solve 3+5x=1

OpenStudy (anonymous):

so when you find the decimal point it would be 0.08 * 100 = 8?

OpenStudy (xapproachesinfinity):

you know what forgot that let's do it the way we do it in real life i paid 17.45$ in total the chair costs 15.99 so the extra tax that i paid is 1.46 now how much percent is 1.46 of 15.99 to find out we just divd 1.46 by 15.99 then we get 0.091 in percentage it will be 9.1% rounding to the nearest 10th percent that is 9%

OpenStudy (xapproachesinfinity):

@kansy yes that how we make it to percentage

OpenStudy (anonymous):

that was the wrong answer dude :(

OpenStudy (anonymous):

for her question

OpenStudy (xapproachesinfinity):

Okay, enlighten me with your knowledge

OpenStudy (xapproachesinfinity):

Alright I'm waiting...

OpenStudy (xapproachesinfinity):

You said I'm wrong! Can you prove me wrong then!

OpenStudy (xapproachesinfinity):

Ate any, I have to go to bed! Think about it! and post what you come up with and will see which of us is wrong lol

OpenStudy (anonymous):

We put 9% on the homework and it said it was wrong ...

OpenStudy (amigatour):

Check this image.

OpenStudy (xapproachesinfinity):

9.10% to nearest tenth error on my part! But the important is you understand how to get the percentage